Clean Corporation manufactures liquid window cleaner. The following information concerns its work in process:

Beginning inventory, 43,100 partially complete gallons.
Transferred out, 209,700 gallons.
Ending inventory (materials are 19 percent complete; conversion costs are 9 percent complete).
Started this month, 238,600 gallons.
Required:
(a) Compute the equivalent units for materials using the weighted-average method.
  Equivalent units
(b) Compute the equivalent units for conversion costs using the weighted-average method.
  Equivalent units
Answer #1

Equivalent unit = Unit transferred out+(Ending work in process*percent completion)

Ending work in process = 43100+238600-209700 = 72000 Units

a) Equivalent unit of material = 209700+(72000*19%) = 223380 Units

b) Equivalent unit of conversion = 209700+(72000*9%) = 216180 Units